OpenStudy (anonymous):

if f(x)= -5x and g(x)= x-4, find: f(x)*g(x) Can you give me steps to solve this?

OpenStudy (anonymous):

\[ f(x)\cdot g(x) = (-5x)\cdot g(x) = (-5x)\cdot (x-4) \]

OpenStudy (anonymous):

Then you distribute.

OpenStudy (anonymous):

f(x)*g(x)=(-5x)(x-4)=-5x^2+20x